# On Robustness of Vision-Language-Action Model against Multi-Modal Perturbations

### v3

@@ Abstract
In Vision–Language–Action (VLA) models, robustness to real-world perturbations is critical for deployment. Existing methods target simple visual disturbances, overlooking the broader multi-modal perturbations that arise in actions, instructions, environments, and observations. Here, we first evaluate the robustness of mainstream VLAs under 17 perturbations across four modalities. We find (1) actions as the most fragile modality, (2) Existing visual-robust VLA do not gain robustness in other modality, and (3) $\pi_{0}$ demonstrates superior robustness with a diffusion-based action head. To build multi-modal robust VLAs, we propose RobustVLA against perturbations in VLA inputs and outputs. For output robustness, we perform offline robust optimization against worst-case action noise that maximizes mismatch in flow matching objective. This can be seen as adversarial training, label smoothing, and outlier penalization. For input robustness, we enforce consistent actions across input variations that preserve task semantics. To account for multiple perturbations, we formulate robustness as a multi-armed bandit problem and apply an upper confidence bound algorithm to automatically identify the most harmful noise. Experiments on LIBERO demonstrate our RobustVLA delivers absolute gains over baselines of 12.6% on the $\pi_{0}$ backbone and 10.4% on the OpenVLA backbone across all 17 perturbations, achieving 50.6x faster inference than existing visual-robust VLAs, and a 10.4% gain under mixed perturbations. Our RobustVLA is particularly effective on real-world FR5 robot with limited demonstrations, showing absolute gains by 65.6% under perturbations of four modalities. Code and demo videos available at https://anonymous.4open.science/r/RobustVLA-283D.
1School of Artificial Intelligence, Beihang University
2School of Computer Science and Engineering, Beihang University
3Institute of Artificial Intelligence, Peking University
4Department of Computer Science and Engineering, The Chinese University of Hong Kong
5Department of Computer Science and Technology, Tsinghua University
6PKU-Psibot Lab
@@ 1 Introduction
To better understand the robustness of VLAs beyond visual uncertainties, as shown in Fig. 1, we evaluate and enhance the robustness of VLAs to multi-modal perturbations. We begin by evaluating the robustness of mainstream VLAs against 17 uncertainties across four modalities. Our findings are threefold: (1) action is the most fragile modality, (2) existing visual-robust VLAs do not show improvements in other modalities, and (3) $\pi_{0}$ (Black et al., 2024) demonstrates superior robustness with a diffusion-based action head and outperforms OpenVLA (Kim et al., 2025) by its large and diverse training corpus. Based on these results, we recommend that robust VLAs focus on robustness in all modalities and build on $\pi_{0}$ backbone as a starting point.
Building on evaluation results, we propose RobustVLA, which handles multi-modal uncertainties in both VLA inputs and outputs. RobustVLA is based on the $\pi_{0}$ backbone and generalizes naturally to OpenVLA. For robustness against VLA output, we derive the worst-case action deviation from the flow-matching objective, then match the action head with both the original and worst-case action distributions. This process can be seen as a combination of flow matching with noisy action distributions, label smoothing, and outlier penalization. For robustness against VLA inputs, we ensure that the noise does not alter the semantics of the current state, so the optimal action remains invariant. We thus regularize the objective to maintain consistent output actions across diverse input perturbations. To balance the various types and modalities of perturbations, we frame the problem as a multi-armed bandit and employ the upper confidence bound (UCB) algorithm (Auer, 2002) to select the most harmful perturbation for training. On LIBERO benchmark, RobustVLA achieves absolute gains of 12.6% on the $\pi_{0}$ backbone and 10.4% on the OpenVLA backbone across 17 perturbations, reaching 50.6x faster inference speed than existing visual-robust VLAs and achieving a 10.4% gain under mixed perturbations. In real-world deployment on the FR5 robot with limited demonstrations, RobustVLA further shows absolute gains of 65.6% under perturbations spanning four modalities.
@@ 2 Related Work
Robust Decision Making. Before the advent of VLAs, robust decision-making was primarily explored within the framework of RL, specifically through robust MDPs (Nilim & El Ghaoui, 2005; Iyengar, 2005). Uncertainties in these settings can arise from various components of MDPs, including environment transitions (Pinto et al., 2017; Mankowitz et al., 2019), actions (Tessler et al., 2019), states (Zhang et al., 2020; 2021), and rewards (Wang et al., 2020). In environments with simulators available, RL agents can learn robust policies through minimax optimization against worst-case adversaries. However, in the case of VLAs, only offline datasets are available, akin to the settings of behavior cloning (Schaal, 1996) and offline RL (Levine et al., 2020). Achieving robust decision-making in the absence of an interactive environment is more challenging, as policies under uncertainty may lead to actions outside the distribution of the original dataset, causing OOD transitions. Consequently, robustness is typically achieved in such settings for states (Shen et al., 2020; Yang et al., 2022; Rigter et al., 2022) and environment transitions (Panaganti et al., 2022; 2023; Seo et al., 2024), with the goal of retaining the original policy despite deviations. However, two major challenges remain when applying these techniques to VLAs. First, it remains unclear how to robustly handle a diverse range of perturbations, with existing methods achieving robustness only against a limited set of environmental uncertainties (Agrawal et al., 2023). Second, it is yet unknown how to achieve action-robust offline RL, as OOD transitions are inevitable in real-world scenarios.
@@ 3.1 Experiment Settings
Evaluated Algorithms. We consider OpenVLA (Kim et al., 2025), $\pi_{0}$-FAST (Pertsch et al., 2025) and $\pi_{0}$ (Black et al., 2024) as representative of autoregressive and diffusion-based VLAs, using their publicly released checkpoint. For robust VLA methods, we consider BYOVLA (Hancock et al., 2025) on $\pi_{0}$ backbone. Due to the lack of publicly available code and insufficient implementation details, we were unable to reproduce GEVRM (Zhang et al., 2025) for fair comparison.
@@ 3.2 Experiment Results
Existing visual-robust VLAs do not show improvements in other modalities. Next, we investigate the robustness of BYOVLA, an off-the-shelf visual-robust VLA that identify and replace the sensitive regions in visual input. As shown in Fig. 3(b), BYOVLA improves performance under Gaussian noise by 7.3% and dead pixels by 22.3%, yet the average gain across all visual corruptions is only 4.0%. Moreover, we observe no measurable improvement ($0.0\%$) in non-visual modalities. These results suggest that existing visual-robust VLAs fall short under multi-modal uncertainties.
$\pi_{0}$ demonstrates superior robustness with a diffusion-based action head. As shown in Fig. 3(c), $\pi_{0}$ outperforms OpenVLA and $\pi_{0}$-FAST by 27.9% and 5.1%, respectively. Since $\pi_{0}$ and $\pi_{0}$-FAST share the same VLM backbone, the performance gap highlights the robustness advantage of the diffusion-based action head over the autoregressive FAST tokenizer. OpenVLA performs even worse, likely due to both its smaller and less diverse training dataset and the limitations of its autoregressive action head. These findings motivate us to focus subsequent robustness-enhancing methods on the $\pi_{0}$ backbone with its diffusion-based action head.
@@ 4 RobustVLA
To counter multi-modal uncertainties, we propose RobustVLA, a framework that enhances robustness against both VLA inputs and outputs. RobustVLA is based on $\pi_{0}$ and naturally extends to OpenVLA.
@@ 4.1 Robustness against VLA Outputs
Ensuring robustness against action output can be hard, since we are equipped with an offline dataset, and any action noises inevitably make subsequent transitions OOD. In our paper, build on $\pi_{0}$ backbone with diffusion-based action head, we first derive $\ell_{p}$ bounded worst-case action noise by maximizing the flow matching loss, then performing robust optimization against such noise.
@@
Worst-Case Action Noise. We define worst-case action noise in VLAs as an $\ell_{p}$-bounded noise that maximally reduces the success rate. In RL, this is typically approximated by a value function, which estimates the quality of the current state-action pair. However, in VLA, estimating action quality is more challenging due to the behavior cloning nature, long horizons, sparse rewards, and complex environmental dynamics. Fortunately, offline demonstrations $(\mathbf{o}_{t},A_{t})$ provide actions that are more likely to lead to success and can be learned through flow matching. Therefore, we use the loss of the flow matching objective, $||v_{\theta}(\hat{A}_{t}^{\tau},\mathbf{o}_{t})-u(\hat{A}_{t}^{\tau}|\hat{A}_{t}^{1})||^{2}$, as an empirical measure of action quality.
@@
Robustness Against Action Noise. With the worst-case action noise $\delta$, enhancing robustness against worst-case $\delta$ requires minimizing the conditional flow matching objective against the worst-case noise. To stabilize the result without noise, we additionally maintain the original $\pi_{0}$ objective, which provides optimal tradeoff between robustness and accuracy (Zhang et al., 2019):
@@
Remark 2. Robustness against action noise can also be interpreted as a form of label smoothing (Müller et al., 2019). Injecting noise into actions makes the learned flow less certain, discouraging overconfident decisions and overfitting to specific actions. This yields better generalization and encourages more stochastic behavior that covers a wider range of plausible actions, a property shown to support robust decision-making in practice (Eysenbach & Levine, 2021).
@@
Generalizing to Autoregressive VLAs. For VLAs that outputs discrete action tokens, take OpenVLA as an example, we perturb the actions before binning to maximize its cross entropy loss, with robust objective following Eqn. 4. The perturbations are constrained so that the action output remains within the original bin and its adjacent bins. This ensures that, even when errors occur, the outputs stay in the neighborhood of the correct action and the worst-case robustness is improved.
@@ 4.2 Robustness against VLA Inputs
where $\lambda_{in}$ balances the $\pi_{0}$ loss and the input robustness term. Finally, to enhance local smoothness of VLAs, we add a $\ell_{p}$ bounded observation noise $\eta$ to maximize the flow matching loss of $\mathcal{L}_{in}^{\tau}$ computed via PGD (Madry et al., 2017), which works well empirically. For autoregressive VLAs, we apply the perturbation to observations and then map to action tokens, the remaining procedure is unchanged.
@@
The pseudocode of our RobustVLA is given in Appendix. D
@@ 5 RobustVLA Experiments
Experiment Setting. Consistent with the robustness evaluation in Section 3.1, we assess all methods under the same 17 perturbations using the LIBERO benchmark and its recommended setup. We conduct our main experiment on $\pi_{0}$ backbone due to its superior robustness compared with OpenVLA. We compare BYOVLA (Hancock et al., 2025), an off-the-shelf robust VLA method designed for visual uncertainties as our baseline. Our approach is termed RobustVLA. For our main experiments, we also consider two ablations of our RobustVLA remove either input or output regularization for better understanding, denoted RobustVLA w/o in and RobustVLA w/o out, respectively. We also evaluate our RobustVLA on the OpenVLA backbone under the same setting. All reported improvements are absolute gains in success rate, expressed in percentage points.
Implementation Details. We follow the default training recipe for $\pi_{0}$ and OpenVLA as recommended in their original codebase. For our RobustVLA, we set $\lambda_{in}$, $\lambda_{out}$ as 1, action noise $\delta$ as 0.03, and observation noise $\eta$ as $8/255$. All baselines use the same set of hyperparameter and implementations. We leave additional implementation details in Appendix. B.1. Code available at https://anonymous.4open.science/r/RobustVLA-283D.
@@ 5.1 RobustVLA on $\pi_{0}$ backbone
RobustVLA is more robust on $\pi_{0}$ backbone. We compare RobustVLA against all non-ablated baselines under 17 perturbations. As shown in Table. 1, RobustVLA gains higher robustness on all 17 perturbations, outperforming $\pi_{0}$ by 14.0% and baseline BYOVLA by 12.6% on average robustness. Under clean conditions without perturbation, RobustVLA remains competitive with $\pi_{0}$ (95.5% vs. 96.0%), showing robustness gains do not come at the expense of clean performance.
@@
Noise Modality Noise type $\pi_{0}$ BYOVLA Ours w/o in Ours w/o out RobustVLA (Ours)
w/o Noise 96.0 95.2 96.0 94.9 95.5
Action Uniform Noise 63.5 62.0 67.3 66.8 69.8
Gaussian Noise 31.4 32.1 36.3 31.9 36.0
Action Bias 23.0 21.2 44.9 37.0 42.3
Random Flips 52.7 51.6 56.9 54.0 58.7
Sudden Spikes 51.7 51.1 58.5 52.7 59.7
Observation Gaussian Noise 51.4 58.7 55.7 94.5 93.8
Dead Pixel 20.8 43.1 40.8 90.8 93.8
Motion Blur 93.7 95.2 94.1 95.0 95.5
Color Jitter 61.7 54.2 53.9 58.7 69.5
Image Rotation 73.3 77.7 64.1 94.0 94.4
Image Shift 74.6 70.4 63.2 89.2 92.7
Environment External Force 37.1 37.3 39.0 37.9 40.8
Irrelevant Objects 93.1 93.7 91.2 89.9 94.2
Lighting Variation 94.3 95.0 94.6 94.3 95.6
Instruction Lexical Transform 78.7 79.5 81.5 77.7 91.3
Syntactic Transform 84.7 85.8 84.0 82.3 93.9
Adversarial Prompts 79.2 80.1 75.7 72.2 80.2
Average 62.6 64.0 64.8 71.7 76.6

### v4

@@ Abstract
In Vision–Language–Action (VLA) models, robustness to real-world perturbations is critical for deployment. Existing methods target simple visual disturbances, overlooking the broader multi-modal perturbations that arise in actions, instructions, environments, and observations. Here, we first evaluate the robustness of mainstream VLAs under 17 perturbations across four modalities. We find (1) actions as the most fragile modality, (2) Existing visual-robust VLA do not gain robustness in other modality, and (3) $\pi_{0}$ demonstrates superior robustness. To build multi-modal robust VLAs, we propose RobustVLA against perturbations in VLA inputs and outputs. For output robustness, we perform offline robust optimization against worst-case action noise that maximizes mismatch in flow matching objective. This can be seen as adversarial training, label smoothing, and outlier penalization. For input robustness, we enforce consistent actions across input variations that preserve task semantics. To account for multiple perturbations, we formulate robustness as a multi-armed bandit problem and apply an upper confidence bound algorithm to automatically identify the most harmful noise. Experiments on LIBERO demonstrate our RobustVLA delivers absolute gains over baselines of 12.6% on the $\pi_{0}$ backbone and 10.4% on the OpenVLA backbone across all 17 perturbations, achieving 50.6x faster inference than existing visual-robust BYOVLA that requires external LLMs, and a 10.4% gain under mixed perturbations. On the real-world FR5 robot, under four types of multimodal perturbations, RobustVLA shows strong low-data performance, outperforming $\pi_{0}$ by $65.6\%$ success rate with 25 demonstrations. Even with abundant demos, our method still outperform $\pi_{0}$ by 30% success rate. Code and demo videos available at https://github.com/gakakulicc/RobustVLA.
1Beihang University, 2Peking University, 3The Chinese University of Hong Kong, 4PKU-Psibot Lab, 5Tsinghua University, 6Zhongguancun Laboratory, 7Hefei Comprehensive National Science Center
@@ 1 Introduction
To better understand the robustness of VLAs beyond visual uncertainties, as shown in Fig. 1, we evaluate and enhance the robustness of VLAs to multi-modal perturbations. We begin by evaluating the robustness of mainstream VLAs against 17 uncertainties across four modalities. Our findings are threefold: (1) action is the most fragile modality, (2) existing visual-robust VLAs do not show improvements in other modalities, and (3) $\pi_{0}$ (Black et al., 2024) demonstrates superior robustness, outperforming OpenVLA (Kim et al., 2025) and $\pi_{0}$-FAST (Pertsch et al., 2025) by large margins. Based on these results, we recommend that robust VLAs focus on robustness in all modalities and build on $\pi_{0}$ backbone as a starting point.
Building on evaluation results, we propose RobustVLA, which handles multi-modal uncertainties in both VLA inputs and outputs. RobustVLA is based on the $\pi_{0}$ backbone and generalizes naturally to other VLAs. For robustness against VLA output, we derive the worst-case action deviation from the flow-matching objective, then match the action head with both the original and worst-case action distributions. This process can be seen as a combination of flow matching with noisy action distributions, label smoothing, and outlier penalization. For robustness against VLA inputs, we ensure that the noise does not alter the semantics of the current state, so the optimal action remains invariant. We thus regularize the objective to maintain consistent output actions across diverse input perturbations. To balance various types of perturbations, we frame the problem as a multi-armed bandit and employ the upper confidence bound (UCB) algorithm (Auer, 2002) to select the most harmful perturbation for training. On LIBERO benchmark, RobustVLA achieves absolute gains of 12.6% on the $\pi_{0}$ backbone and 10.4% on the OpenVLA backbone across 17 perturbations, being 50.6x faster than visual-robust BYOVLA that requires external LLMs, and achieving a 10.4% gain under mixed perturbations. In real-world deployment with four multimodal perturbations, RobustVLA performs strongly in the low-data regime, outperforming $\pi_{0}$ by $65.6\%$ with 25 demos, and still achieves $30\%$ higher success rate than $\pi_{0}$ with 100 demos, where the performance of $\pi_{0}$ saturates.
@@ 2 Related Work
Robust Decision Making. Before the advent of VLAs, robust decision-making was primarily explored within the framework of RL, specifically through robust MDPs (Nilim and El Ghaoui, 2005; Iyengar, 2005). Uncertainties in these settings can arise from various components of MDPs, including environment transitions (Pinto et al., 2017; Mankowitz et al., 2019), actions (Tessler et al., 2019), states (Zhang et al., 2020; 2021), and rewards (Wang et al., 2020). In environments with simulators available, RL agents can learn robust policies through minimax optimization against worst-case adversaries. However, in the case of VLAs, only offline datasets are available, akin to the settings of behavior cloning (Schaal, 1996) and offline RL (Levine et al., 2020). Achieving robust decision-making in the absence of an interactive environment is more challenging, as policies under uncertainty may lead to actions outside the distribution of the original dataset, causing OOD transitions. Consequently, robustness is typically achieved in such settings for states (Shen et al., 2020; Yang et al., 2022; Rigter et al., 2022) and environment transitions (Panaganti et al., 2022; 2023; Seo et al., 2024), with the goal of retaining the original policy despite deviations. However, two major challenges remain when applying these techniques to VLAs. First, it remains unclear how to robustly handle a diverse range of perturbations, with existing methods achieving robustness only against a limited set of environmental uncertainties (Agrawal et al., 2023). Second, it is yet unknown how to achieve action-robust offline RL, as OOD transitions are inevitable in real-world scenarios.
@@ 3.1 Experiment Settings
Evaluated Algorithms. We consider OpenVLA (Kim et al., 2025), $\pi_{0}$-FAST (Pertsch et al., 2025) and $\pi_{0}$ (Black et al., 2024) as representative of autoregressive and diffusion-based VLAs, using their publicly released checkpoint. For robust VLA methods, we consider BYOVLA (Hancock et al., 2025) on $\pi_{0}$ backbone , we use the official implementation and apply its robustness enhancement procedure only to the visual modality. Due to the lack of publicly available code and insufficient implementation details, we were unable to reproduce GEVRM (Zhang et al., 2025) for fair comparison.
@@ 3.2 Experiment Results
Existing visual-robust VLAs do not show improvements in other modalities. Next, we investigate the robustness of BYOVLA, an off-the-shelf visual-robust VLA that identify and replace the sensitive regions in visual input. As shown in Fig. 3(b), BYOVLA improves performance under Gaussian noise by 7.3% and dead pixels by 22.3%, yet the average gain across all visual corruptions is only 4.0%. Moreover, we observe no measurable improvement ($0.0\%$) in non-visual modalities. These results indicate that visual robustness alone does not translate into broader robustness across other modalities, motivating our study of comprehensive multi-modal robustness in VLA models.
$\pi_{0}$ demonstrates higher robustness than OpenVLA and $\pi_{0}$-FAST. As shown in Fig. 3(c), $\pi_{0}$ outperforms OpenVLA and $\pi_{0}$-FAST by 27.9% and 5.1%, respectively. Since $\pi_{0}$ and $\pi_{0}$-FAST share the same VLM backbone, this difference suggests an advantage of the diffusion-based action head. We evaluate both models under varying noise and observe that the robustness of $\pi_{0}$-FAST degrades faster than $\pi_{0}$ (Appendix C.2), which motivates us to develop RobustVLA on the $\pi_{0}$ backbone.
@@ 4 RobustVLA
To counter multi-modal uncertainties, we propose RobustVLA, a fine-tuning framework to enhance robustness against uncertainties in VLA inputs and outputs. We take $\pi_{0}$ with rectified flow matching as an example due to its popularity, while our method naturally extends to VLAs with general diffusion-based action head and autoregressive VLAs like OpenVLA.
@@ 4.1 Robustness against VLA Outputs
Ensuring robustness against action output can be hard, since we are equipped with an offline dataset, and any action noises inevitably make subsequent transitions OOD, which amplifies subsequent errors (Levine et al., 2020). In our paper, taking $\pi_{0}$ with a rectified flow matching action head, we first derive $\ell_{p}$ bounded worst-case action noise by maximizing the flow matching loss, then performing robust optimization against such noise.
@@
Worst-Case Action Noise. We define worst-case action noise in VLAs as an $\ell_{p}$-bounded noise that maximally reduces the success rate. In robust RL literature, since success rate is hard to measure per step, researchers assume existing well-trained policy have high chance to succeed, and maximizing deviation to current policy leads to successful attacks (Huang et al., 2017). In VLAs, offline demonstrations $(\mathbf{o}_{t},A_{t})$ provide actions that are more likely to lead to success and can be learned through flow matching. Therefore, we use the loss of the flow matching objective, $||v_{\theta}(\hat{A}_{t}^{\tau},\mathbf{o}_{t})-u(\hat{A}_{t}^{\tau}|\hat{A}_{t}^{1})||^{2}$, as an empirical measure of action quality. We also conduct a pilot study in Appendix. C.5, which empirically shows flow matching loss is strongly correlated with success rate, with significant Pearson correlation of r=0.95r=-0.95, $p<0.05$.
@@
Robustness Against Action Noise. With the worst-case action noise $\delta$, enhancing robustness against worst-case $\delta$ requires minimizing the conditional flow matching objective against the worst-case noise. To stabilize the result without noise, we use the TRADES objective that maintain the original $\pi_{0}$ loss, which provides optimal tradeoff between robustness and accuracy (Zhang et al., 2019):
@@
Remark 2. Robustness against action noise can also be interpreted as a form of label smoothing (Müller et al., 2019). Injecting noise into actions makes the learned flow less certain, discouraging overconfident decisions and overfitting to specific actions. This yields better generalization and encourages more stochastic behavior that covers a wider range of plausible actions, a property shown to support robust decision-making in practice (Eysenbach and Levine, 2021).
@@
Generalizing to Other VLAs. For general diffusion-based VLAs, just ignore the assumption of $u(\hat{A}_{t}^{\tau}|\hat{A}_{t}^{1})=u(A_{t}^{\tau}|A_{t}^{1})-\delta$ of rectified flow and use the first line of Eqn. 3 to generate the worst-case action perturbation. For autoregressive VLAs, take OpenVLA as an example, we perturb the actions before binning to maximize its cross entropy loss, with robust objective following Eqn. 4. The perturbations are constrained so that the action output remains within the original bin and its adjacent bins. This ensures that, even when errors occur, the outputs stay in the neighborhood of the correct action and the worst-case robustness is improved.
@@ 4.2 Robustness against VLA Inputs
where $\lambda_{in}$ balances the $\pi_{0}$ loss and the input robustness term. Finally, to enhance local smoothness of VLAs, we add a $\ell_{p}$ bounded observation noise $\eta$ to maximize the flow matching loss of $\mathcal{L}_{in}^{\tau}$ computed via PGD (Madry et al., 2017), which works well empirically.
Generalizing to Other VLAs. For other diffusion-based VLAs, our method can be applied by changing the rectified flow matching loss directly. For autoregressive VLAs, we apply the perturbation to observations and then map to action tokens, the remaining procedure is unchanged.
@@
The pseudocode of our RobustVLA is given in Appendix. D.
@@ 5 RobustVLA Experiments
Experiment Setting. Consistent with the robustness evaluation in Section 3.1, we assess all methods under the same 17 perturbations using the LIBERO benchmark and its recommended setup. We conduct our main experiment on $\pi_{0}$ backbone due to its superior robustness compared with OpenVLA. We compare BYOVLA (Hancock et al., 2025), an off-the-shelf robust VLA method designed for visual uncertainties as our baseline. Our approach is termed RobustVLA. For our main experiments, we also consider four ablations of our RobustVLA, including: (1) a domain randomization (DR) variant that trains $\pi_{0}$ using randomly selected VLA input perturbations, (2) our RobustVLA without input regularization (w/o in), (3) our RobustVLA without output regularization (w/o out), (4) our RobustVLA without UCB for balancing multiple input noises (w/o UCB). We also evaluate our RobustVLA on the OpenVLA backbone under the same setting. All reported improvements are absolute gains in success rate, expressed in percentage points.
Implementation Details. We follow the default training recipe for $\pi_{0}$ and OpenVLA as recommended in their original codebase. For our RobustVLA, we set $\lambda_{in}$, $\lambda_{out}$ as 1, action noise $\delta$ as 0.03, and observation noise $\eta$ as $8/255$. All baselines use the same set of hyperparameter and implementations. See implementation details and additional training details in Appendix B.
@@ 5.1 RobustVLA on $\pi_{0}$ backbone
RobustVLA is more robust on $\pi_{0}$ backbone. We compare RobustVLA against all non-ablated baselines under 17 perturbations. As shown in Table. 1, RobustVLA gains higher robustness on all 17 perturbations, outperforming $\pi_{0}$ by 14.0% and baseline BYOVLA by 12.6% on average robustness. Under clean conditions without perturbation, RobustVLA remains competitive with $\pi_{0}$ (95.5% vs. 96.0%), showing robustness gains do not come at the expense of clean performance. Notably, our method also generalizes to unseen perturbations not encountered during training, such as external forces, demonstrating robustness beyond the set of uncertainties considered during training. A separate analysis on LIBERO-long highlights the potential of RobustVLA on complex and long-horizon tasks, outperforming $\pi_{0}$ by 19.61% in success rate. See results in Appendix. C.3. Results also confirmed action as the modality worth further investigation, with our method achieving +8.0% success rate on average, comparing with +17.2% increase in uncertainties applied to VLA input.
@@
For DR and Ours w/o UCB, several patterns emerge. First, DR performs poorly. While DR improves robustness to some observation noises, it fails under most environment and instruction noises, likely because it overfits to a small subset of easy perturbations. Second, Ours w/o UCB shows a 7.3% drop in robustness, indicating that UCB is crucial for balancing multiple perturbations rather than letting the model overfit to a single dominant noise source. Comparing with Ours w/o in that trains on output noise only, adding exposure to diverse input noises adds further robustness gains, even without UCB.
@@
Noise Modality Noise type $\pi_{0}$ DR BYOVLA w/o in w/o out w/o UCB RobustVLA(Ours)
w/o Noise 96.0 94.8 95.2 96.0 94.9 94.9 95.5
Action Uniform Noise 63.5 61.2 62.0 67.3 66.8 69.5 69.8
Gaussian Noise 31.4 30.1 32.1 36.3 31.9 35.9 36.0
Action Bias 23.0 27.6 21.2 44.9 37.0 41.3 42.3
Random Flips 52.7 48.5 51.6 56.9 54.0 57.6 58.7
Sudden Spikes 51.7 54.2 51.1 58.5 52.7 62.3 59.7
Observation Gaussian Noise 51.4 64.1 58.7 55.7 94.5 84.3 93.8
Dead Pixel 20.8 54.5 43.1 40.8 90.8 78.5 93.8
Motion Blur 93.7 88.7 95.2 94.1 95.0 95.2 95.5
Color Jitter 61.7 54.0 54.2 53.9 58.7 62.3 69.5
Image Rotation 73.3 85.9 77.7 64.1 94.0 70.7 94.4
Image Shift 74.6 72.3 70.4 63.2 89.2 62.3 92.7
Environment External Force 37.1 31.9 37.3 39.0 37.9 39.0 40.8
Irrelevant Objects 93.1 80.8 93.7 91.2 89.9 91.2 94.2
Lighting Variation 94.3 89.0 95.0 94.6 94.3 92.4 95.6
Instruction Lexical Transform 78.7 78.8 79.5 81.5 77.7 77.1 91.3
Syntactic Transform 84.7 72.3 85.8 84.0 82.3 83.6 93.9
Adversarial Prompts 79.2 56.7 80.1 75.7 72.2 74.6 80.2
Average 62.6 61.8 64.0 64.8 71.7 69.3 76.6